www.ably.io
Back

Incident log archive

All times are shown in UTC

August 2017

23rd August 2017 03:35:00 PM

Elevated error rates

15:35 UTC A faulty deploy in ap-southeast-2 is causing higher than normal error rates worldwide. We are reverting the deploy.

16:23 UTC Error rates are back to normal in all regions.

Resolved

1 minute
22nd August 2017 04:14:14 AM

Our automated health check system has reported an issue with realtime cluster health in eu-central-1-a

This incident was created automatically by our automated health check system as it has identified a fault. We are now looking into this issue.

22nd Aug 04:14 AM

Our health check system has reported this issue as resolved.
We will continue to investigate the issue and will update this incident shortly.

22nd Aug 06:59 AM

There was a momentary loss of connectivity in the EU Central region. It lasted for approximately 1-2 minutes and resolved itself.

Resolved

less than a minute

July 2017

23rd July 2017 11:00:00 AM

High latencies and error rates in EU West

We are seeing intermittent high latencies being reported in our EU West datacenter.

We are investigating the root cause of the issue now. In the mean time, we are adding more capacity to the region.

23rd Jul 02:40 PM

We identified the root cause of the issue was an internal monitoring system that went rogue and was generating significant load in the cluster.

We have disabled that specific monitoring system and will roll out a fix before turning that system back on.

All services are now back to normal.

Resolved

5 minutes
18th July 2017 06:08:01 PM

Higher latencies in US-East-1

We are seeing growing latencies in our US East (North Virginia) data center with latencies being reported up to twice as high as normal.

This latency increase is linked to a recent deploy today.

We are investigating the root cause and will soon bring on additional capacity to mitigate the issue whilst we resolve the underlying problem.

18th Jul 09:21 PM

At 18:40 the latencies returned to normal in us-east-1 following additional capacity being introduced into the cluster.

18th Jul 09:22 PM

The issue was resolved by rolling back to a previous version of the realtime system. We will be trying to reproduce this issue in our lab environments now before rolling this out in production again.

Resolved

1 minute
5th July 2017 02:27:00 PM

Performance issues

From 14:27 UTC we are seeing performance issue across the cluster on some apps and channels, and are investigating

Update: The incident was resolved 14:51 UTC. The cause was a configuration issue combined with a transient inter-region netsplit.

Resolved

3 minutes

June 2017

2nd June 2017 04:34:00 PM

Performance issues in us-east

We are experiencing performance issues in us-east-1. We are investigating. All other regions are unaffected.

Edit: advisory expanded to all regions

2nd Jun 05:14 PM

The issue is now resolved.

We identified the root cause as memory starvation on an instance, however this in itself should have caused little or even no disruption whilst our health services recovered the problem.

The continued high error rates were unfortunately down to a problem in the Gossip protocol. We had indefinitely pending nodes that were not there i.e. they had been disposed of by the health services, yet the nodes in the cluster still reported the node as there. So the root cause of this issue is a Gossip fault which we are investigating now.

Resolved

less than a minute

May 2017

31st May 2017 03:03:13 PM

us-east-1 traffic spike

We are seeing a tremendous spike in traffic arriving at us-east-1. Some of those requests are being re-routed automatically to other regions.

We're initially adding more capacity to address the issue whilst we investigate.

31st May 03:17 PM

The traffic spike subsided within 5 minutes. We'll ensure there is a generous amount of extra capacity in place in this region for 24 hours whilst we investigate where this traffic originated.

Resolved

1 minute
23rd May 2017 06:00:16 PM

us-east-1 slow requests

We are seeing a higher rate of slow requests in our US East 1 datacenter. Until we understand the underlying cause, we are temporarily adding more capacity to fix the problem.

23rd May 10:33 PM

We identified a number of nodes with memory exhaustion. We fixed the issue quickly but and are now investigating the root cause of the issue.

Resolved

less than a minute
22nd May 2017 08:52:04 AM

Minor network disruption in us-west-1

Our external monitoring systems reported a very brief period of network connectivity issues in our US West 1 datacenter. The issue resolved itself within a few minutes.

Resolved

less than a minute
16th May 2017 10:19:42 PM

Abrupt exits in Asia causing disruption primarily in Asia

A set of nodes in Asia have exited abruptly causing significant disruptions in Asia to channels. We are also seeing a knock-on effect in other regions due to roles constantly migrating around the cluster to compensate for the abruptly exiting nodes.

We are investigating and attempting to resolve the abrupt exits in Asia first so that a normal service can be restored.

We will then review the root cause.

16th May 10:41 PM

We have restored stability in Asia, however we are seeing gossip instability in other regions. We continue to investigate the root cause, but our priority remains to restore stability in other regions. We are recycling unstable nodes in other regions to resolve the issue.

16th May 11:03 PM

The global realtime service stability has been resumed in full.

We are now investigating the root cause of the gossip instability to understand why a consensus was not reached soon after the abrupt exits in Asia. We'll also been investigating the cause of the abrupt exits we saw in Asia.

Once our investigation is complete, we'll be writing up a post mortem on the issues this evening.

17th May 02:01 PM

Upon inspection of the logs at the time of the instability, it's clear we have a bug in our Gossip service that resulted in ghosts remaining in the ring that were in fact gone. We are investigating a few avenues which should lead to fixes soon.

Resolved

about 15 hours
10th May 2017 11:25:00 AM

Further performance issues and timeouts

We have been seeing relatively high rates of timeout errors in all datacenters starting at 11:25 UTC; we are investigating

10th May 01:54 PM

The error rates have returned to normal. This was due to a new protocol deploy that should have been compatible, but unfortunately introduced unexpected communication issues.

Resolved

1 minute
10th May 2017 01:12:00 AM

Performance issues and timeouts

A number of nodes in the system at 1:12am became unstable on our gossip ring and were unable to resolve the issues automatically. This has resulted in an increased number of timeout errors due to some nodes having a different view of the gossip ring than the remaining nodes in the system.

We are looking into the issue.

10th May 08:22 AM

As of 08:18 UTC, the issue has been resolved in all datacenters apart from eu-central-1. We are still investigating higher than normal error rates in eu-central-1.

10th May 10:01 AM

The issues in eu-central are now resolved in full. We will provide a full update on the root cause an action we've taken soon.

Resolved

25 minutes

April 2017

6th April 2017 10:07:22 PM

eu-west-1-a spike

Our automated systems detected a spike in an eu-west-1 datacentre that was abrupt, resulted in very slow responses, and was resolved within around 4 minutes.

Resolved

about 1 month
6th April 2017 08:39:07 AM

Abrupt instance crash & ring instability

We have been notified of an abrupt instance crash in Singapore. Normally this has no side effects other than a small delay whilst all nodes agree that the node has gone, and as our design ensures that data is stored in at least two locations, no data is lost.

However, we are seeing some cluster instability in that the nodes appear to be unable to agree with each other on which nodes are active or not.

We are investigating the root cause and can see error rates have increased somewhat (1-2%) for some customers.

6th Apr 09:14 AM

We redeployed the instance which were unable to come to consensus on the ring. This resolved the issue.

We have however raised an internal ticket to now investigate why the ring was unable to reach stability within at most 10-15s, which is what we expect when a node is abruptly terminated.

Resolved

about 1 month
3rd April 2017 12:56:10 PM

Cluster partitioning

We are seeing cluster instability reported again, following a 10 minute period this morning where this happened too.

We are investigating further why this has happened twice now in one day.

3rd Apr 01:09 PM

It appears that the instability was caused by genuine network partitioning issues again.

We will continue to keep an eye on network issues.

Resolved

about 1 month